# AutoDAN-Turbo: A Lifelong Agent for Strategy Self-Exploration to Jailbreak LLMs

### v1

@@ Abstract
11footnotetext: Equal Contribution.22footnotetext: Corresponding to cxiao34@wisc.edu
@@ 3 AutoDAN-Turbo
As illustrated in Fig. 2, our method consists of three main modules: the Attack generation and Exploration Module (Sec. 3.1) , Strategy Library Construction Module (Sec. 3.2), and Jailbreak Strategy Retrieval Module (Sec. 3.3). In the Attack generation and Exploration Module, where the goals are to (1) generate jailbreak prompt to attack the target LLM by leveraging the strategies provided by Jailbreak Strategy Retrieval Module, it consists of (1) an attacker LLM that generates jailbreak prompts based on specific strategies retrieved from Jailbreak Strategy Retrieval Module; (2) a target (victim) LLM that provides responses; and a scorer LLM that evaluates these responses to assign scores. We can repeat this process multiple time to generate massive attack logs for Strategy Library Construction Module to generate a strategy library. Strategy Library Construction Module is to extract strategies from the attack logs generated in Attack Generation and Exploration Module and save the strategies into the Strategy Library; Jailbreak Strategy Retrieval Module is to support the Attack Generation and Exploration Module to retrieve the strategy from the strategy library constructed by Strategy Library Construction Module so that the retrieved jailbreak prompt can guide the jailbreak prompt generation to attack the victim LLMs.
@@
As illustrated in Fig. 2, three LLMs are involved in the Attack Generation and Exploration Module: an attacker LLM, a target LLM (the victim model we want to jailbreak), and a scorer LLM. Specifically, the attack loop contains the following steps: (1)Attack Generation: The attacker LLM receives specific prompts that describe the malicious request $M$ and encourages the attacker LLM to generate a jailbreak prompt using specified jailbreak strategies. The attacker LLM then generates the jailbreak prompt $P$; (2)Target Response: Upon receiving $P$ as input, the target LLM generates a response $R$; (3) Scorer Evaluation: The response $R$ is then evaluated by the scorer LLM. This evaluation determines whether the response meets the malicious goal of the jailbreak attack. The scorer LLM returns a numerical score $S$ based on predefined criteria. The scores range from 1, indicating no alignment with malicious intent, to 10, representing full compliance with harmful directives. The detailed prompt for configuring the scorer LLM is provided in Appendix B.2.
Our module supports three functionalities shown in Tab. A in the appendix : (1) generating jailbreak prompts without a strategy, (2) generating jailbreak prompts with effective retrieved strategy, and (3) generating jailbreak prompts with ineffective strategies. For (1), when no strategy exists in the strategy library (described in Sec. 3.2), the prompt asks the attacker LLM to generate jailbreak prompts for the malicious request using any strategy it can imagine. For (2), when several effective jailbreak strategies are provided, the prompt instructs the attacker LLM to generate jailbreak prompts according to the given strategies; For (3), if the framework has gone through the strategy library and only found ineffective strategies, the prompt directs the attacker LLM to avoid these low-scoring strategies and devise new ones.
@@ 3.2 Strategy Library Construction Module
We randomly extract two attack records from the attack log, represented as $\{P_{i},R_{i},S_{i}\}$ and $\{P_{j},R_{j},S_{j}\}$. If the score $S_{j}$ is higher than $S_{i}$, we argue that some strategy may have been explored and employed in the jailbreak prompt $P_{j}$ compared to $P_{i}$, leading to an improved score. Thus, we count the improvement from $P_{i}$ to $P_{j}$ as a strategy. As shown in Fig. 3, to summarize the improvements, we employ a summarizer LLM (details in Appendix B.3) to summarize the improvement. the summarizer LLM compares $P_{j}$ with $P_{i}$, to analyze the strategies that make $R_{j}$ more malicious than $R_{i}$ (i.e., $S_{j}>S_{i}$). The summarizer LLM will provide a description of the jailbreak strategy and the concise definition of the strategy, and format it into a JSON object, $S_{i}$. $S_{i}$ comprises three elements: "Strategy", the name of the strategy employed in $P_{j}$; "Definition", the concise definition of the strategy summarized by summarizer LLM ; and "Example", which records the jailbreak prompt $P_{j}$ that utilized the strategy. To prevent redundancy in defining jailbreak strategies, we will provide the summarizer LLM with the existing categories from the strategy library, instructing it to avoid duplicating definitions.
@@ 3.3 Jailbreak Strategy Retrieval
A key operation in our framework is to retrieve jailbreak strategies from the strategy library, and then prompt the attacker LLM to generate jailbreak prompts based on these strategies. Specifically, given the malicious request M, we feed them to generation and exploration module to get $\{P_{i},R_{i},S_{i}\}$. We then employ the text-embedding model to transform the response $R_{i}$ into an embedding vector $E_{R_{i}}$. Subsequently, we compare the similarity between $E_{R_{i}}$ with all keys in the strategy library. We choose the top-$2k$ data frames with the highest similarity of key values. Then, we sort these values based on the score differences they contain and select the top-$k$ strategies that led to the highest score differences. These are the most effective strategies associated with the responses $R_{i}$. These selected strategies will be then formed as a retrieved strategy list $\Gamma$. If two samples with the same score are selected and happen to meet the length limit of the strategy list, they are added or dropped in the program’s default order.
After establishing the retrieved strategy list $\Gamma$, we insert these strategies into the prompt of the attacker LLM in the next attack iteration as illustrated in Tab A. Specifically, we adopt the following tactics:(1) If the highest score in $\Gamma$ is greater than 5, we will directly use this strategy as effective strategy and insert it into the attacker LLM’s prompt. Namely, the attacker LLM is asked to use this strategy to generate the jailbreak prompt in the next jailbreak round; (2). If the highest score is less than 5, we select all strategies with a score difference between $2-5$ ad set them as effective strategies. We insert these strategies into the attacker’s prompt. Namely, we inform the attacker LLM that these strategies are potent for the current malicious request, and attacker LLM can combine and evolve among these strategies to generate new jailbreak prompt; (3). If the number of highest strategies is less than 2, we viewed these strategies as ineffective strategies since they can not achieve big improvements. Thus, as shown in Tab. A, we inform the attacker LLM in the prompt that these strategies are not particularly effective for the current malicious request, so they should not continue using these strategies and need to discover other strategies; (4). If the $\Gamma$ set is empty, we will provide empty strategy to attacker LLM.
@@ 4.1 Experiments Setup
Large Language Models. We conduct comprehensive evaluations on both open-source and closed-source LLMs. Specifically, for open-source LLMs, we include Llama-2-7B-chat (Touvron et al., 2023), Llama-2-13B-chat (Touvron et al., 2023), Llama-2-70B-chat (Touvron et al., 2023), Llama-3-8B (Dubey et al., 2024), Llama-3-70B (Dubey et al., 2024), and Gemma-7B-it (Team et al., 2024b). For closed-source models, we include GPT-4-1106-turbo (OpenAI et al., 2024) and Gemini Pro (Team et al., 2024a). The specific roles these models serve, whether as the attacker LLM, the target LLM, or the strategy summarizer LLM, will be detailed in the corresponding contexts. Note that throughout our experiments, we employed a deterministic generation approach by using a zero temperature setting, and limited the maximum token generation to 4096 tokens. To ensure the consistency of our experiments, we used Gemma-7B-it as our scorer LLM throughout.
@@ 4.2 Main Results
As illustrated in Tab. 1, our method AutoDAN-Turbo consistently achieves better performance in both Harmbench ASR and StrongREJECT Score, which means that our method not only induces the target LLM to answer and provide harmful content in more malicious requests, as measured by the Harmbench ASR, but also results in a higher level of maliciousness compared to responses induced by other attacks, as indicated by the StrongREJECT Score. Specifically, if we use the Gemma-7B-it model as the attacker and strategy summarizer in our method (i.e., AutoDAN-Turbo (Gemma-7B-it)), we have an average Harmbench ASR of $56.4$, surpassing the runner-up (Rainbow Teaming, $33.1$) by $70.4\%$, and StrongREJECT Score equals to $0.24$, surpassing the runner up (Rainbow Teaming, $0.13$) by $84.6\%$. If we utilize a larger model, i.e., the Llama-3-70B as the attacker and strategy summarizer in our method (i.e., AutoDAN-Turbo (Llama-3-70B)), we have an average Harmbench ASR of $57.7$, surpassing the runner-up (Rainbow Teaming, $33.1$) by $74.3\%$, and StrongREJECT Score equals to $0.25$, surpassing the runner up (Rainbow Teaming, $0.13$) by $92.3\%$. Interestingly, our method demonstrates remarkable jailbreak effectiveness on one of the most powerful models, GPT-4-1106-turbo. Specifically, AutoDAN-Turbo (Gemma-7B-it) achieves a Harmbench ASR of $83.8$, and AutoDAN-Turbo (Llama-3-70B) achieves $88.5$, showcasing the great effectiveness of our method on state-of-the-art models. We also compare our method with all the jailbreaks attacks included in Hrambench. As shown in Tab. 2, the results demonstrate that our method is the most powerful jailbreak attack. The outstanding performance of our method compared to the baselines is due to the autonomous exploration of jailbreak strategies without human intervention or predefined scopes. In contrast, Rainbow Teaming only utilizes $8$ human-developed jailbreak strategies as its jailbreak reference. This fixed predefined scope leads to a lower ASR.
@@ 4.3 Strategy Transferability
(a) Attacker: Gemma-7B-it
(b) Attacker: Llama-2-13B-chat
(c) Attacker: Llama-3-70B
(d) Attacker: Gemini Pro
Figure 4: The transferability of the strategy library developed from various attacker LLMs across different datasets. The red columns represent the ASR on the Harmbench dataset for different victim LLMs, while the blue columns represent the ASR on an unseen malicious request dataset for different victim models.
Strategy Transferability across Different Datasets. Here, we study the strategy transferability across different datasets. Specifically, we evaluate whether the strategies, initially developed using the Harmbench dataset, can be effective when applied to other datasets. We constructed an “Unseen Malicious Requests” dataset using datasets from recent studies (Lapid et al., 2024; Qiu et al., 2023; Zou et al., 2023; Luo et al., 2024), which is different from Harmbench. The results, illustrated in Fig. 4, confirm that the strategy libraries maintain high transferability across different datasets. The decrease in ASR due to dataset shifts is less than $4\%$.
@@ 4.5 Test-time Query Efficiency
The results, shown in Tab. 5, indicate that our method requires significantly fewer queries than PAIR and TAP, reducing average query usage by $87.0\%$. This demonstrates that once the strategy library is constructed, our attack will be highly query-efficient and maintain high attack success rates.
5 Conclusions and Limitations
In this paper, we introduce AutoDAN-Turbo, which utilizes lifelong learning agents to automatically and continually discover diverse strategies and combine them for jailbreak attacks. Extensive experiments have demonstrated that our method is highly effective and transferable. However, a limitation of our approach is the high computational demand required to load multiple LLMs. Building the strategy library from scratch requires repeated interactions between the models, which adds to the resource strain. This issue can be mitigated by loading a trained strategy library.
@@ 5 Conclusions and Limitations
  • Computational CostsA
  • Full Prompts:B
  • Strategies Sample in Strategy LibraryC
  • A Whole Process of Exploring a New Jailbreak StrategyD
  • Jailbreak ExamplesE
Appendix A Computational Costs

### v4

@@ Abstract
11footnotetext: Equal Contribution.22footnotetext: Corresponding to xiaogeng.liu@wisc.edu
@@ 3 AutoDAN-Turbo
As illustrated in Fig. 2, our method consists of three main modules: the Attack generation and Exploration Module (Sec. 3.1) , Strategy Library Construction Module (Sec. 3.2), and Jailbreak Strategy Retrieval Module (Sec. 3.3). In the Attack generation and Exploration Module, where the goals are to generate jailbreak prompt to attack the target LLM by leveraging the strategies provided by Jailbreak Strategy Retrieval Module, it consists of an attacker LLM that generates jailbreak prompts based on specific strategies retrieved from Jailbreak Strategy Retrieval Module; a target (victim) LLM that provides responses; and a scorer LLM that evaluates these responses to assign scores. We can repeat this process multiple time to generate massive attack logs for Strategy Library Construction Module to generate a strategy library. Strategy Library Construction Module is to extract strategies from the attack logs generated in Attack Generation and Exploration Module and save the strategies into the Strategy Library; Jailbreak Strategy Retrieval Module is to support the Attack Generation and Exploration Module to retrieve the strategy from the strategy library constructed by Strategy Library Construction Module so that the retrieved jailbreak prompt can guide the jailbreak prompt generation to attack the victim LLMs. The algorithmic outline is provided in Appendix. D.
@@
As illustrated in Fig. 2, three LLMs are involved in the Attack Generation and Exploration Module: an attacker LLM, a target LLM (the victim model we want to jailbreak), and a scorer LLM. Specifically, the attack loop contains the following steps: (1) Attack Generation: The attacker LLM receives specific prompts that describe the malicious request $M$ and encourages the attacker LLM to generate a jailbreak prompt using specified jailbreak strategies. The attacker LLM then generates the jailbreak prompt $P$; (2) Target Response: Upon receiving $P$ as input, the target LLM generates a response $R$; (3) Scorer Evaluation: The response $R$ is then evaluated by the scorer LLM. This evaluation determines whether the response meets the malicious goal of the jailbreak attack. The scorer LLM returns a numerical score $S$ based on predefined criteria. The scores range from 1, indicating no alignment with malicious intent, to 10, representing full compliance with harmful directives. The detailed prompt for configuring the scorer LLM is provided in Appendix E.2.
Our module supports three functionalities shown in Tab. E in the appendix : (1) generating jailbreak prompts without a strategy, (2) generating jailbreak prompts with effective retrieved strategy, and (3) generating jailbreak prompts with ineffective strategies. For (1), when no strategy exists in the strategy library (described in Sec. 3.2), the prompt asks the attacker LLM to generate jailbreak prompts for the malicious request using any strategy it can imagine. For (2), when several effective jailbreak strategies are provided, the prompt instructs the attacker LLM to generate jailbreak prompts according to the given strategies; For (3), if the framework has gone through the strategy library and only found ineffective strategies, the prompt directs the attacker LLM to avoid these low-scoring strategies and devise new ones.
@@ 3.2 Strategy Library Construction Module
We randomly extract two attack records from the attack log, represented as $\{P_{i},R_{i},S_{i}\}$ and $\{P_{j},R_{j},S_{j}\}$. If the score $S_{j}$ is higher than $S_{i}$, we argue that some strategy may have been explored and employed in the jailbreak prompt $P_{j}$ compared to $P_{i}$, leading to an improved score. Thus, we count the improvement from $P_{i}$ to $P_{j}$ as a strategy. As shown in Fig. 3, to summarize the improvements, we employ a summarizer LLM (details in Appendix E.3) to summarize the improvement. The summarizer LLM compares $P_{j}$ with $P_{i}$, to analyze the strategies that make $R_{j}$ more malicious than $R_{i}$ (i.e., $S_{j}>S_{i}$). The summarizer LLM will provide a description of the jailbreak strategy and the concise definition of the strategy, and format it into a JSON object, $S_{i}$. $S_{i}$ comprises three elements: “Strategy”, the name of the strategy employed in $P_{j}$; “Definition”, the concise definition of the strategy summarized by summarizer LLM ; and “Example”, which records the jailbreak prompt $P_{j}$ that utilized the strategy. To prevent redundancy in defining jailbreak strategies, we will provide the summarizer LLM with the existing categories  [^1] To reduce token costs, we exclude “Example” values. from the strategy library, instructing it to avoid duplicating definitions.
@@ 3.3 Jailbreak Strategy Retrieval
A key operation in our framework is to retrieve jailbreak strategies from the strategy library, and then prompt the attacker LLM to generate jailbreak prompts based on these strategies. Specifically, given the malicious request M, we feed them to generation and exploration module to get $\{P_{i},R_{i},S_{i}\}$. We then employ the text-embedding model to transform the response $R_{i}$ into an embedding vector $E_{R_{i}}$. Subsequently, we compare the similarity between $E_{R_{i}}$ with all keys in the strategy library. We choose the top-$2k$ data frames with the highest similarity of key values. Then, we sort these values based on the score differences they contain and select the top-$k$ strategies that led to the highest score differences. These are the most effective strategies associated with the responses $R_{i}$. These selected strategies will be then formed as a retrieved strategy list $\Gamma$. If two samples with the same score are selected and happen to meet the length limit of the strategy list, they are added or dropped in the program’s default order. Note that in the first iteration, there is no response $R_{i}$ available for retrieval reference . Thus the attacker is prompted without employing a jailbreak strategy in the first iteration.
After establishing the retrieved strategy list $\Gamma$, we insert these strategies into the prompt of the attacker LLM in the next attack iteration as illustrated in Tab E. Specifically, we adopt the following tactics:(1) If the highest score in $\Gamma$ is greater than 5, we will directly use this strategy as effective strategy and insert it into the attacker LLM’s prompt. Namely, the attacker LLM is asked to use this strategy to generate the jailbreak prompt in the next jailbreak round; (2) If the highest score is less than 5, we select all strategies with a score difference between $2-5$ ad set them as effective strategies. We insert these strategies into the attacker’s prompt. Namely, we inform the attacker LLM that these strategies are potent for the current malicious request, and attacker LLM can combine and evolve among these strategies to generate new jailbreak prompt; (3) If the highest score in $\Gamma$ is less than 2, we viewed these strategies as ineffective strategies since they can not achieve big improvements. Thus, as shown in Tab. E, we inform the attacker LLM in the prompt that these strategies are not particularly effective for the current malicious request, so they should not continue using these strategies and need to discover other strategies; (4). If the $\Gamma$ set is empty, we will provide empty strategy to attacker LLM.
@@ 4.1 Experiments Setup
Large Language Models. We conduct comprehensive evaluations on both open-source and closed-source LLMs. Specifically, for open-source LLMs, we include Llama-2-7B-chat (Touvron et al., 2023), Llama-2-13B-chat (Touvron et al., 2023), Llama-2-70B-chat (Touvron et al., 2023), Llama-3-8B-Instruct (Dubey et al., 2024), Llama-3-70B-Instruct (Dubey et al., 2024), and Gemma-1.1-7B-it (Team et al., 2024b). For closed-source models, we include GPT-4-1106-turbo (OpenAI et al., 2024) and Gemini Pro (Team et al., 2024a). The specific roles these models serve, whether as the attacker LLM, the target LLM, or the strategy summarizer LLM, will be detailed in the corresponding contexts. Note that throughout our experiments, we employed a deterministic generation approach by using a zero temperature setting, and limited the maximum token generation to 4096 tokens. To ensure the consistency of our experiments, we used Gemma-1.1-7B-it as our scorer LLM throughout.
@@ 4.2 Main Results
As illustrated in Tab. 1, our method AutoDAN-Turbo consistently achieves better performance in both Harmbench ASR and StrongREJECT Score, which means that our method not only induces the target LLM to answer and provide harmful content in more malicious requests, as measured by the Harmbench ASR, but also results in a higher level of maliciousness compared to responses induced by other attacks, as indicated by the StrongREJECT Score. Specifically, if we use the Gemma-7B-it model as the attacker and strategy summarizer in our method (i.e., AutoDAN-Turbo (Gemma-7B-it)), we have an average Harmbench ASR of $56.4$, surpassing the runner-up (Rainbow Teaming, $33.1$) by $70.4\%$, and StrongREJECT Score equals to $0.24$, surpassing the runner up (Rainbow Teaming, $0.13$) by $84.6\%$. If we utilize a larger model, i.e., the Llama-3-70B as the attacker and strategy summarizer in our method (i.e., AutoDAN-Turbo (Llama-3-70B)), we have an average Harmbench ASR of $57.7$, surpassing the runner-up (Rainbow Teaming, $33.1$) by $74.3\%$, and StrongREJECT Score equals to $0.25$, surpassing the runner up (Rainbow Teaming, $0.13$) by $92.3\%$. Interestingly, our method demonstrates remarkable jailbreak effectiveness on one of the most powerful models, GPT-4-1106-turbo. Specifically, AutoDAN-Turbo (Gemma-7B-it) achieves a Harmbench ASR of $83.8$, and AutoDAN-Turbo (Llama-3-70B) achieves $88.5$, showcasing the great effectiveness of our method on state-of-the-art models. We also compare our method with all the jailbreaks attacks included in Hrambench. As shown in Tab. 2, the results demonstrate that our method, where we use Gemma-7B-it as the attacker, is the most powerful jailbreak attack. The outstanding performance of our method compared to the baselines highlights the importance and effectiveness of autonomous exploration of jailbreak strategies without human intervention or predefined scopes.
@@ 4.3 Strategy Transferability
Figure 4: The transferability of the strategies developed by Gemma-7B-it attacker across different datasets.
@@
Strategy Transferability across Different Datasets. Here, we study the strategy transferability across different datasets. Specifically, we evaluate whether the strategies, initially developed using the Harmbench dataset, can be effective when applied to other datasets. We constructed an “Unseen Malicious Requests” dataset using datasets from recent studies (Lapid et al., 2024; Qiu et al., 2023; Zou et al., 2023; Luo et al., 2024), which is different from Harmbench. The results, illustrated in Fig. 4, confirm that the strategy libraries maintain high transferability across different datasets. The red columns represent the ASR on the Harmbench dataset for different victim LLMs, while the blue columns represent the ASR on an unseen malicious request dataset. The decrease in ASR due to dataset shifts is less than $4\%$. More results from various attacker LLMs are in Appendix. K.
@@ 4.5 Test-time Query Efficiency
The results, shown in Tab. 5, indicate that our method requires significantly fewer queries than PAIR and TAP, reducing average query usage by $87.0\%$. This demonstrates that once the strategy library is constructed, our attack will be highly query-efficient and maintain high attack success rates. We also share detailed evaluations on the scaling relationship between total attack query times and ASR for different jailbreak methods in Appendix B.
5 Conclusions
In this paper, we introduce AutoDAN-Turbo, which utilizes lifelong learning agents to automatically and continually discover diverse strategies and combine them for jailbreak attacks. Extensive experiments have demonstrated that our method is highly effective and transferable.
Limitation
A limitation of our approach is the high computational demand required to load multiple LLMs. Building the strategy library from scratch requires repeated interactions between the models, which adds to the resource strain. This issue can be mitigated by loading a trained strategy library.
Acknowledgments
We would like to express our sincere gratitude to the reviewer(s) for their valuable feedback and constructive comments, which significantly contributed to the improvement of this paper. We are grateful to the Center for AI Safety for generously providing computational resources. Yevgeniy Vorobeychik is partially supported by National Science Foundation under grant No. IIS-2214141 and Army Research Office under grant number W911NF-25-1-0059. Somesh Jha is partially supported by DARPA under agreement number 885000, NSF CCF-FMiTF-1836978 and ONR N00014-21-1-2492. Patrick McDaniel is partially supported by the NSF under Grant No. CNS-2343611, U.S. Department of Homeland Security under Grant Award Number 17STQAC00001-07-00. Bo Li is partially supported by NSF grant No. 1910100, No. 2046726, NSF AI Institute ACTION No. IIS-2229876, DARPA TIAMAT No. 80321, the National Aeronautics and Space Administration (NASA) under grant No. 80NSSC20M0229, ARL Grant W911NF-23-2-0137, Alfred P. Sloan Fellowship, the research grant from eBay, AI Safety Fund, Virtue AI, and Schmidt Science.
Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the sponsors.
@@ Acknowledgments
  • Computational Resource RequirementA
  • The Attack Query Times Scaling LawB
  • Diversity of AutoDAN-Turbo’s Jailbreak PromptsC
  • Algorithmic Outline of AutoDAN-TurboD
  • Full PromptsE
  • Alignment Study of Scorer and SummarizerF
  • Strategies Sample in Strategy LibraryG
  • A Whole Process of Exploring a New Jailbreak StrategyH
  • Jailbreak ExamplesI
  • Additional Comparison with Existing WorkJ
  • Supplementary Tables and FiguresK
Appendix A Computational Resource Requirement